OpenStudy (anonymous):

Verify the identity. cot(x-pi/2 = -tan x I have one explaination...but it doesnt particularly PROVE

OpenStudy (anonymous):

I feel like this is too litterary and not enough MATH basis... The original cofunction identity read: cot(pi/2-x)=tan x But by "flipping" or inverting what is in the parenthesis, you have now inverted tan, forming -tan. This is in conjunction with the Odd/Even Identities.

OpenStudy (anonymous):

im looking for a more numbery....kind of way of putting it...or well you understnad

OpenStudy (jdoe0001):

$$\bf cot\left(x-\cfrac{\pi}{2}\right) \implies \cfrac{ cos\left(x-\cfrac{\pi}{2}\right) }{ sin\left(x-\cfrac{\pi}{2}\right) }\\ \cfrac{ cos(x)cos\left(\cfrac{\pi}{2}\right)+sin(x)sin\left(\cfrac{\pi}{2}\right) }{ sin(x)cos\left(\cfrac{\pi}{2}\right)-cos(x)sin\left(\cfrac{\pi}{2}\right) } $$

OpenStudy (jdoe0001):

now you'd just need to keep in mind what is the \(\bf sin\left(\cfrac{\pi}{2}\right)\)? what is the \(\bf cos\left(\cfrac{\pi}{2}\right)\)?

OpenStudy (anonymous):

sin pi/2 is cospi/2 and inverted for cos...

OpenStudy (anonymous):

i cant use pictures in my explaination @cwrw238 ...

OpenStudy (anonymous):

@jdoe0001 ..was i correct?

OpenStudy (jdoe0001):

well, yes look at your unit circle, sine of pi/2 and cos of pi/2 are pretty conspicuous

OpenStudy (anonymous):

haha yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(0) and 1

OpenStudy (anonymous):

well 1 and 0

OpenStudy (jdoe0001):

anything there, multiplying for \(\bf cos\left(\cfrac{\pi}{2}\right)\) will be multiplied by 0 and 0 times anything = 0 anything there multiplying for \(\bf sin\left(\cfrac{\pi}{2}\right)\) will be multiplied by 1, and 1 times anything is "anything"

OpenStudy (anonymous):

ok..wait a second i need to catch up a little ok?

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

so you get one over one?...

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

sinx/-cosx

OpenStudy (anonymous):

OOOH

OpenStudy (jdoe0001):

\(\bf \cfrac{ cos(x)cos\left(\cfrac{\pi}{2}\right)+sin(x)sin\left(\cfrac{\pi}{2}\right) }{ sin(x)cos\left(\cfrac{\pi}{2}\right)-cos(x)sin\left(\cfrac{\pi}{2}\right) }\\ \cfrac{ 0+sin(x)(1) }{ 0-cos(x)(-1) }\)

OpenStudy (anonymous):

THANK YOUUU

OpenStudy (anonymous):

I see now :3

OpenStudy (anonymous):

fanks

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

woops I even put a -1 for ... anyhow \(\bf \cfrac{ 0+sin(x)(1) }{ 0-cos(x)(1) }\) better

OpenStudy (anonymous):

haha I understand :)

OpenStudy (jdoe0001):

just had a typo

OpenStudy (anonymous):

thanks a heap. im glad you explainedstep by step!

OpenStudy (jdoe0001):

np